OpenStudy (anonymous):

(1 - 7x)(1 + 9x) need help solving useing the foil method can anyone help me plz i have to have this done today

hero (hero):

(a + b)(a + b) = a(a + b) + b(a + b) = a^2 + ab + ab + b^2 = a^2 + 2ab + b^2

hero (hero):

(2x + 4) (x + 5) = 2x(x + 5) + 4(x + 5) = 2x^2 + 10x + 4x + 25 = 2x^2 + 14x + 25

OpenStudy (anonymous):

that is so confusing

OpenStudy (anonymous):

thats not even my problem

hero (hero):

I know. It's an example

hero (hero):

Which is exactly what you need. You need to do the actual problem yourself

OpenStudy (texaschic101):

FOIL method..... First, outside, inside, last.... (1 - 7x)(1 + 9x) Lets take the First numbers...and multiply... 1 x 1 = 1 now take the outside numbers ... 1 x 9x = 9x now take the inside numbers... (7x) x 1 = 7x now take the last numbers... (7x) x 9x = 63x^2 Now lets put it all together.... 1 + 9x + 7x + 63x^2 (combine like terms) 1 + 16x + 63x^2 lets write it like this... 63x^2 + 16x + 1

OpenStudy (anonymous):

i have examples but i dont under stand them im doing online work for school so i need help going through it step by step its not easy trying to do all this on your own when u havnt realy done it before

hero (hero):

Have you seen this before? a(b + c) = ab + ac That's distributive rule

OpenStudy (texaschic101):

what problems do u need help with

hero (hero):

Distributive Rule and FOIL are the same method. The only difference is, distributive rule is an actual property of real numbers. FOIL is an acronym.

OpenStudy (anonymous):

the next prob is worse then this one (3m^3 -1/2 y)(3m^3 -1/2 y)

hero (hero):

Not really. It's the same method, just with higher degree terms.

OpenStudy (anonymous):

but see i dont understand it at all can u help me sep by step to solve this i have no teacher and theres no one here who understands

hero (hero):

There's another method you might understand.

hero (hero):